# Hier-SLAM: Scaling-up Semantics in SLAM with a Hierarchically Categorical Gaussian Splatting

### v3

@@ Abstract
We propose Hier-SLAM, a semantic 3D Gaussian Splatting SLAM method featuring a novel hierarchical categorical representation, which enables accurate global 3D semantic mapping, scaling-up capability, and explicit semantic label prediction in the 3D world. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making it particularly challenging and costly for scene understanding. To address this problem, we introduce a novel hierarchical representation that encodes semantic information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs). We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Furthermore, we enhance the whole SLAM system, resulting in improved tracking and mapping performance. Our Hier-SLAM outperforms existing dense SLAM methods in both mapping and tracking accuracy, while achieving a 2x operation speed-up. Additionally, it exhibits competitive performance in rendering semantic segmentation in small synthetic scenes, with significantly reduced storage and training time requirements. Rendering FPS impressively reaches 2,000 with semantic information and 3,000 without it. Most notably, it showcases the capability of handling the complex real-world scene with more than 500 semantic classes, highlighting its valuable scaling-up capability.
@@ I INTRODUCTION
Figure 1: (a). The global 3D Gaussian map generated by Hier-SLAM with learned semantic labels is shown on the left. The hierarchical structure of the semantic information is organized on the right, considering both semantic and geometric attributes (blue box). The proposed hierarchical categorical representation compresses semantic data, reducing both memory usage and training time of the semantic SLAM. (b). The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
@@
Thus, it is highly desirable to extend the original 3D Gaussian Splatting with semantic capabilities while preserving its advantageous probabilistic representation. A straightforward approach would be to augment 3D points with a discrete semantic label and parameterize its distribution with a categorical discrete distribution, i.e., a flat Softmax embedding representation. However, 3D Gaussian Splatting is already a storage-intensive representation [14, 15], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would result in significantly increased storage demands and processing time, growing linearly with the number of semantic classes. This makes it particularly impractical for complex scene understanding. Recent works formulate semantic classes using non-distributional approaches to handle this complexity. The work [16] directly learns a 3-channel RGB visualization for semantic maps instead of the semantic label learning. Another work [17] uses a flat semantic representation with supervision from pre-trained foundation models, yet it can only produces a 3D semantic feature map, rather than an explicit 3D semantic label map.
Unlike flat representations, semantic information naturally organises into a hierarchical structure of classes, as illustrated in Fig. 1. This hierarchical relationship can be effectively represented as a tree structure, allowing for efficient encoding of extensive information with a relatively small number of nodes, i.e., a compact code. For instance, a binary tree with a depth of $10$ can cover $2^{10}$ classes, enabling the representation of $1,024$ classes using just 20 codes (i.e., $2\times 10$, through 2-dimensional Softmax coding for each level).
@@
3)
We conduct experiments on both synthetic and real-world datasets. The results demonstrate that our SLAM system outperforms existing methods in localization and mapping performance while achieving faster speeds. In small synthetic scenes, our method achieves competitive performance in rendering semantic segmentation. In complex real-world scenes, our approach, for the first time, demonstrates a valuable scaling-up capability, successfully handling more than 500 semantic classes—an important step toward the semantic understanding of complex environments.
@@ II RELATED WORK
3D Gaussian Splatting has emerged as a promising 3D representation recently. With the usage of 3D Gaussian Splatting, SplaTAM [9] leverages silhouette guidance for pose estimation and map reconstruction in RGBD SLAM systems. Gaussian Splatting SLAM [10] implements both monocular and RGBD SLAM using 3D Gaussian Splatting. 3D Gaussian Splatting has demonstrated its strong capabilities across various Gaussian Splatting SLAM tasks [11, 12, 13]. However, integrating semantic understanding into SLAM tasks makes optimization particularly challenging, as it combines three high-dimensional optimization problems with different value ranges and convergence characteristics that to be optimized jointly. In this paper, we leverage hierarchical coding for semantic information and employ a suitable optimization strategy to ensure effective optimization.
@@
Semantic SLAM has been a longstanding research topic in the field of computer vision and robotics [4, 18, 19, 20, 21]. Many works [5, 22, 23] have utilized the neural implicit representation for semantic mapping and localization tasks. DNS-SLAM [5] leverages 2D semantic priors combined with a coarse-to-fine geometry representation to integrate semantic information into the established map. SNI-SLAM [22] incorporates appearance, geometry, and semantic features into a collaborative feature space to enhance the robustness of the entire SLAM system. However, these methods are constrained by the limitations of neural implicit map representations, which is known to suffer from slow convergence, which leads to inefficiency and performance degradation when combined with semantic objectives [24, 25]. In contrast, Gaussian Splatting offers advantages with its fast rendering performance and high-density reconstruction quality.
@@
With the recent emergence of 3D Gaussian Splatting, the work [16] integrated additional RGB 3-channels to learn semantic visualization map, rather than true semantic understanding. SemGauss-SLAM [17] employs a flat semantic representation, supervised by a large pre-trained foundation model. However, these methods neglect the natural hierarchical characteristics of the real world. Furthermore, the reliance on large foundation models increases the complexity of the neural network and its computational demands, with performance heavily dependent on the embeddings from these pre-trained models. In this paper, we introduce a simple yet effective hierarchical representation for semantic understanding, eliminating the dependency on foundation models, enabling a coarse-to-fine semantic understanding for the unknown environments.
@@ III-A Hierarchical representation
Tree Parametrization. We propose a hierarchical tree representation to encode semantic information, represented as $G=(V,E)$. The node set $V=\cup_{l=0}^{L}\left\{v_{l}\right\}$ comprises all classes, where $\left\{v_{l}\right\}$ represents the set of nodes at the $l$-th level of the tree. The edge set $E=\cup_{m=0}^{M}\left\{e_{m}\right\}$ captures the subordination relationships, encompassing both semantic attribution and geometric prior knowledge. Similarly we use the subscript $m$ to indicate the level of the tree. In this way, the $i$-th semantic class $g^{i}$, regarded as a single leaf node in the tree view, can be expressed hierarchically as:
$g_{leaf}^{i}=\{v_{l}^{i},e_{m}^{i}\mid l=0,1,...,L;\;m=0,1,...,M\},$
@@
which corresponds to the root-to-leaf path: $g_{leaf}^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{M-2}}v_{L-1}^{i}\xrightarrow{e_{M}}v_{L}^{i}$. Take a leaf node class ’Wall’ as an example, a 4-level tree coding can be as: $\left\{v_{0}^{wall}:\texttt{Background}\right\}\rightarrow\left\{v_{1}^{wall}:\texttt{Structure}\right\}\rightarrow\left\{v_{2}^{wall}:\texttt{Plane}\right\}\rightarrow\left\{v_{3}^{wall}:\texttt{Wall}\right\}$. Among these nodes, the relationships such as ‘include’ and ‘possessing’ are represented by the edge information $e_{m}:\rightarrow$. In this way, any semantic concept can be coded in a progressive, hierarchical manner, incorporating both semantic and geometric perspectives. Moreover, the standard flat representation can be seen as a single-level tree coding from the hierarchical viewpoint.
@@
LLM-based Tree Generation. We utilize Large Language Models (LLMs), specifically GPT-4o-mini [26], to generate the hierarchical tree representation due to its efficient performance. Specifically, the set of original semantic classes are input into the LLMs to cluster them into coarser-level classes. This process is repeated layer by layer from leaf to root, ultimately forming the complete hierarchical tree. However, when dealing with a large number of classes in a complex environment, the results are often unsatisfactory because the LLM tends to cluster only a subset of the input classes, leaving out many classes and incorrectly including numerous unseen classes in the hierarchy.
@@
To address this issue, we employ a loop-based critic operation, including an LLM followed by a validator. Specifically, during the clustering process from the $l$-level to the $(l-1)$-level, the $l$-level semantic classes $\left\{v_{l}\right\}$ are used as the initial prompt input to the LLM, awaiting clustering. The LLM then generates the clustering result $\left\{v_{l}^{\prime}\right\}\rightarrow\left\{v^{\prime}_{l-1}\right\}$. By comparing the LLM’s results and the prompt input, the validator will identify three components: the successfully grouped nodes $\left\{v_{l}^{\prime}\right\}^{\text{success}}$, the unseen classes $\left\{v_{l}^{\prime}\right\}^{\text{unseen}}$, and the omitted semantic nodes $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}$. The successfully grouped nodes $\left\{v_{l}^{\prime}\right\}^{\text{success}}$ will be retained, while the unseen classes $\left\{v_{l}^{\prime}\right\}^{\text{unseen}}$ will be removed. Next, the omitted nodes $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}$ are used as the input prompt for the LLM to do the clustering in the subsequent iteration. Additionally, the clustering nodes $\left\{v^{\prime}_{l-1}\right\}$ generated in the previous iteration are also provided to the LLM as a reference, suggesting that $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}$ can either be clustered into the previously generated clusters or form new groups. This procedure loops until $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}=\emptyset$, indicating that no classes have been omitted. In this way, we obtain all the clustering results from the $l$-level to the $(l-1)$-level. The proposed loop-based critic operation progresses from the leaf-to-root levels, ending when fewer than $\theta$ clusters can be generated by the LLM, where we set $\theta=4$. Worth mentioning is that the tree generation is performed offline before the SLAM operation.
@@
$\bm{h}=f(\bm{h}^{l})\in\mathbb{R}^{N},\quad\bm{h}^{l}\in\mathbb{R}^{n^{l}},\quad l=0,1,...,L$ (1)
where we use $l$ to represents $l$-th level of the tree and $f$ stands for the concatenation operation. As shown in Fig. 2, the overall dimension of the hierarchical embedding is the sum of the dimensions across all levels $N=\sum_{l=0}^{L}n^{l}$, where the dimension $n_{l}$ of each embedding $\bm{h}^{l}$ depends on the maximum number of nodes at the $l$-th level.
@@ III-B Hierarchical loss
where $\mathcal{P}$ represents the semantic ground truth.
@@ III-C Gaussian Splatting Semantic Mapping and Tracking
In contrast to previous work [9], which employs separate forward and backward rendering modules for different parameters, we adopt unified forward and backward modules that handle all parameters, including semantic, color, depth, and silhouette images.
@@ IV-A Experiment settings
The experiments are conducted on both synthetic and real-world datasets, including 6 scenes from ScanNet [28] and 8 sequences from Replica [27]. Following the evaluation metrics used in previous SLAM works [9, 29], we leverage ATE RMSE (cm) to assess SLAM tracking accuracy. For mapping performance, we use Depth L1 (cm) to evaluate accuracy. To assess image rendering quality, we adopt PSNR (dB), SSIM, and LPIPS metrics. Similar to previous methods [17, 5, 22], due to the lack of direct metrics for evaluating 3D semantic understanding in 3D Gaussian Splatting representations, we rely on 2D semantic segmentation performance, measured by mIoU (mean Intersection over Union across all classes), to reflect global semantic information. To demonstrate the improved efficiency, we also measure the running time of the proposed SLAM method. We compare our method against state-of-the-art dense visual SLAM approaches, including both NeRF-based and 3D Gaussian SLAM methods, to highlight its effectiveness. Additionally, we include state-of-the-art semantic SLAM techniques, covering both NeRF-based and Gaussian-based methods, to showcase our hierarchical semantic understanding and scaling-up capability. The experiments are conducted in the Nvidia L40S GPU. For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly. We set semantic optimization loss weights $\omega_{1}$ and $\omega_{2}$ to1.01.01.01.0 and $0.0$, respectively, for the first $\eta$ iterations, where $\eta$ is set to $15$. Afterwards, $\omega_{1}$ and $\omega_{2}$ are adjusted to $1.0$ and $5.0$, respectively. This means that we first use the Inter-level loss to initialize the hierarchical coding, followed by incorporating the Cross-level loss to refine the embedding. For tracking loss, we set $\delta=0.99$, $w_{1}=1.0$, $w_{2}=0.5$, respectively. For mapping, we set $w_{3}=1.0$, $w_{4}=0.5$, $w_{5}=0.2$, respectively.

### v4

@@ Abstract
We propose Hier-SLAM, a semantic 3D Gaussian Splatting SLAM method featuring a novel hierarchical categorical representation, which enables accurate global 3D semantic mapping, scaling-up capability, and explicit semantic label prediction in the 3D world. The parameter usage in semantic SLAM systems increases significantly with the growing complexity of the environment, making it particularly challenging and costly for scene understanding. To address this problem, we introduce a novel hierarchical representation that encodes semantic information in a compact form into 3D Gaussian Splatting, leveraging the capabilities of large language models (LLMs). We further introduce a novel semantic loss designed to optimize hierarchical semantic information through both inter-level and cross-level optimization. Furthermore, we enhance the whole SLAM system, resulting in improved tracking and mapping performance. Our Hier-SLAM outperforms existing dense SLAM methods in both mapping and tracking accuracy, while achieving a 2x operation speed-up. Additionally, it achieves on-par semantic rendering performance compared to existing methods while significantly reducing storage and training time requirements. Rendering FPS impressively reaches 2,000 with semantic information and 3,000 without it. Most notably, it showcases the capability of handling the complex real-world scene with more than 500 semantic classes, highlighting its valuable scaling-up capability. The open-source code is available at https://github.com/LeeBY68/Hier-SLAM.
@@ I INTRODUCTION
Figure 1: (a). The global 3D Gaussian map generated by Hier-SLAM with learned semantic labels is shown on the left. The hierarchical structure of the semantic information is organized on the right, considering both semantic and geometric attributes (the second blue box). The proposed hierarchical categorical representation compresses semantic data, reducing both memory usage and training time of the semantic SLAM. (b). The rendered semantic map at different levels shows a coarse-to-fine understanding, beneficial for real-world scenarios with shifting perspectives from distant to close.
@@
Thus, it is highly desirable to extend the original 3D Gaussian Splatting with semantic capabilities while preserving its advantageous probabilistic representation. A straightforward approach would be to augment 3D points with a discrete semantic label and parameterize its distribution with a categorical discrete distribution, i.e., a flat Softmax embedding representation. However, 3D Gaussian Splatting is already a storage-intensive representation [14, 15], requiring a large number of 3D primitives with multiple parameters to achieve realistic rendering. Adding semantic distribution parameters would result in significantly increased storage demands and processing time, growing linearly with the number of semantic classes. This makes it particularly impractical for complex scene understanding. Recent works formulate semantic classes using non-distributional approaches to handle this complexity. The work [16] directly learns a 3-channel RGB visualization for semantic maps instead of the semantic label learning. Another work [17] uses a flat semantic representation with supervision from pre-trained foundation models to produces a 3D semantic embedding feature map.
Unlike flat representations, semantic information naturally organises into a hierarchical structure of classes, as illustrated in Fig. 1. This hierarchical relationship can be effectively represented as a tree structure, allowing for efficient encoding of extensive information with a relatively small number of nodes, i.e., a compact code. For instance, a binary tree with a depth of $10$ can cover $2^{10}$ classes, enabling the representation of $1,024$ classes using just symbolic 20 codes (i.e., $2\times 10$, through 2-dimensional Softmax coding for each level).
@@
3) We conduct experiments on both synthetic and real-world datasets. The results demonstrate that our SLAM system outperforms existing methods in localization and mapping performance while achieving faster speeds. For semantic understanding, our method achieves on-par semantic rendering performance while significantly reducing storage and training time. In complex real-world scenes, our approach, for the first time, demonstrates a valuable scaling-up capability, successfully handling more than 500 semantic classes—an important step toward the semantic understanding of complex environments.
@@ II RELATED WORK
3D Gaussian Splatting has emerged as a promising 3D representation recently. With the usage of 3D Gaussian Splatting, SplaTAM [9] leverages silhouette guidance for pose estimation and map reconstruction in RGBD SLAM systems. MonoGS [10] implements both monocular and RGBD SLAM using 3D Gaussian Splatting. 3D Gaussian Splatting has demonstrated its strong capabilities across various Gaussian Splatting SLAM tasks [11, 12, 13]. However, integrating semantic understanding into SLAM tasks makes optimization particularly challenging, as it combines three high-dimensional optimization problems with different value ranges and convergence characteristics that to be optimized jointly. In this paper, we leverage hierarchical coding for semantic information and employ a suitable optimization strategy to ensure effective optimization across the hierarchical representation.
@@
Semantic SLAM has been a longstanding research topic in the field of computer vision and robotics [4, 18, 19, 20, 21]. Many works [5, 22, 23] have utilized the neural implicit representation for semantic mapping and localization tasks. DNS-SLAM [5] leverages 2D semantic priors combined with a coarse-to-fine geometry representation to integrate semantic information into the established map. SNI-SLAM [22] incorporates appearance, geometry, and semantic features into a collaborative feature space to enhance the robustness of the entire SLAM system. However, these methods are constrained by the limitations of neural implicit map representations, which is known to suffer from slow convergence, which leads to inefficiency and performance degradation when combined with semantic objectives [24, 25]. In contrast, Gaussian Splatting offers advantages with its fast rendering performance and high-density reconstruction quality at the same time.
@@
With the recent emergence of 3D Gaussian Splatting, SGS-SLAM [16] integrated additional RGB 3-channels to learn semantic visualization map, rather than true semantic understanding. SemGauss-SLAM [17] employs a flat semantic representation, supervised by a large pre-trained foundation model. However, these methods neglect the natural hierarchical characteristics of the real world. Furthermore, the reliance on large foundation models increases the complexity of the neural network and its computational demands, with performance heavily dependent on the embeddings from these pre-trained models. In this paper, we introduce a simple yet effective hierarchical representation for semantic understanding, eliminating the dependency on foundation models, enabling a coarse-to-fine semantic understanding for the unknown environments.
@@ III-A Hierarchical representation
Tree Parametrization. We propose a hierarchical tree representation to encode semantic information, represented as $G=(V,E)$. The node set $V=\cup_{l=0}^{L}\left\{v_{l}\right\}$ comprises all classes, where $\left\{v_{l}\right\}$ represents the set of nodes at the $l$-th level of the tree. The edge set $E=\cup_{m=0}^{L-1}\left\{e_{m}\right\}$ captures the subordination relationships, encompassing both semantic attribution and geometric prior knowledge. Similarly we use the subscript $m$ to indicate the level of the tree. In this way, the $i$-th semantic class $g^{i}$, regarded as a single leaf node in the tree view, can be expressed hierarchically as:
$g^{i}=\{v_{l}^{i},e_{m}^{i}\mid l=0,1,...,L;\;m=0,1,...,L-1\},$ (1)
@@
which corresponds to the root-to-leaf path: $g^{i}=v_{0}^{i}\xrightarrow{e_{0}}v_{1}^{i}\xrightarrow{e_{1}}\cdots\xrightarrow{e_{L-2}}v_{L-1}^{i}\xrightarrow{e_{L-1}}v_{L}^{i}$. Take a leaf node class ’Wall’ as an example, a 4-level tree coding can be as: $\left\{v_{0}^{\text{wall}}:\texttt{Background}\right\}\rightarrow\left\{v_{1}^{\text{wall}}:\texttt{Structure}\right\}\rightarrow\left\{v_{2}^{\text{wall}}:\texttt{Plane}\right\}\rightarrow\left\{v_{3}^{\text{wall}}:\texttt{Wall}\right\}$. Among these nodes, the relationships such as ‘include’ and ‘possessing’ are represented by the edge information $e_{m}:\rightarrow$. In this way, any semantic concept can be coded in a progressive, hierarchical manner, incorporating both semantic and geometric perspectives. Moreover, the standard flat representation can be seen as a single-level tree coding from the hierarchical viewpoint.
@@
LLM-based Tree Generation. We utilize Large Language Models (LLMs), GPT-4o-mini [26], to generate the hierarchical tree representation due to its efficient performance. Specifically, a set of semantic class labels is provided to the LLMs, which iteratively clusters them into groups, regarding as the coarser-level classes. This process is repeated layer by layer from leaf to root, ultimately forming a complete hierarchical tree. However, when dealing with a large number of semantic class labels in a complex environment, the results are often unsatisfactory because the LLM tends to cluster only a subset of the input classes, leaving out many classes and incorrectly including unseen classes in the hierarchy.
@@
To address this issue, we employ a loop-based critic operation, including an LLM followed by a validator. Specifically, during the clustering process from the $l$-level to the $(l-1)$-level, the $l$-level semantic classes $\left\{v_{l}\right\}$ are used as the prompt input to the LLM. The LLM then generates the clustering result $\left\{v_{l}^{\prime}\right\}\rightarrow\left\{v^{\prime}_{l-1}\right\}$. It is important to note that the clustering result $\left\{v_{l}^{\prime}\right\}$ may differ from the input $\left\{v_{l}\right\}$, as the LLM may introduce unseen classes or omit certain semantic labels. By comparing the clustering result $\left\{v_{l}^{\prime}\right\}$ and the prompt input $\left\{v_{l}\right\}$, the validator will identify three components: the successfully grouped nodes $\left\{v_{l}^{\prime}\right\}^{\text{success}}$, the unseen classes $\left\{v_{l}^{\prime}\right\}^{\text{unseen}}$, and the omitted semantic nodes $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}$. The successfully grouped nodes $\left\{v_{l}^{\prime}\right\}^{\text{success}}$ will be retained, while the unseen classes $\left\{v_{l}^{\prime}\right\}^{\text{unseen}}$ will be removed. Next, the omitted nodes $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}$ are used as the input prompt for the LLM to do the clustering in the subsequent iteration. At the same iteration, the clustering nodes $\left\{v^{\prime}_{l-1}\right\}$ generated by previous iteration are also provided to the LLM as a reference, suggesting that $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}$ can either be clustered into the previously generated clusters or form new groups. This procedure loops until $\left\{v_{l}^{\prime}\right\}^{\text{omitted}}=\emptyset$, indicating that no classes are omitted. In this way, we obtain all clustering results from the $l$-level to the $(l-1)$-level. The proposed loop-based critic operation progresses from leaf to root, terminating when the LLM generates fewer than $\theta$ clusters. We set $\theta=4$, ensuring that the number of nodes in the finest level remains small. It is worth noting that the tree generation is performed offline before the SLAM operation.
@@
$\bm{h}=f(\bm{h}^{l})\in\mathbb{R}^{N},\quad\bm{h}^{l}\in\mathbb{R}^{n},\quad l=0,1,...,L$ (2)
where we use $l$ to represents $l$-th level of the tree and $f$ stands for the concatenation operation. As shown in Fig. 2, the overall dimension of the hierarchical embedding is the sum of the dimensions across all levels $N=\sum_{l=0}^{L}n$, where the dimension $n$ of each embedding $\bm{h}^{l}$ is equal to the maximum number of nodes at the $l$-th level.
@@ III-B Hierarchical loss
where $\mathcal{P}$ denotes semantic ground truth in flat representation.
@@ III-C Gaussian Splatting Semantic Mapping and Tracking
In contrast to previous work [9], which employs separate forward and backward rendering modules for different parameters, we adopt unified forward and backward modules that handle all parameters, including semantic, color, depth, and silhouette images, significantly improving the overall efficiency of the SLAM system.
@@ IV-A Experiment settings
The experiments are conducted on both synthetic and real-world datasets, including 6 scenes from ScanNet [28] and 8 sequences from Replica [27]. Following the evaluation metrics used in previous SLAM works [9, 29], we leverage ATE RMSE (cm) to assess SLAM tracking accuracy. For mapping performance, we use Depth L1 (cm) to evaluate accuracy. To assess image rendering quality, we adopt PSNR (dB), SSIM, and LPIPS metrics. Similar to previous methods [17, 5, 22], due to the lack of direct metrics for evaluating 3D semantic understanding in 3D Gaussian Splatting representations, we rely on 2D semantic segmentation performance, measured by mIoU (mean Intersection over Union across all classes), to reflect global semantic information. To demonstrate the improved efficiency, we also measure the running time of the proposed SLAM method. We compare our method against state-of-the-art dense visual SLAM approaches, including both NeRF-based and 3D Gaussian SLAM methods, to highlight its effectiveness. Additionally, we include state-of-the-art semantic SLAM techniques, covering both NeRF-based and Gaussian-based methods, to showcase our hierarchical semantic understanding and scaling-up capability. The experiments are conducted in the Nvidia L40S GPU. For experimental settings, the semantic embedding of each Gaussian primitive is initialized randomly. We set semantic optimization loss weights $\omega_{1}$ and $\omega_{2}$ to1.01.01.01.0 and $0.0$, respectively, for the first $\eta$ iterations, where $\eta$ is set to $15$. Afterwards, $\omega_{1}$ and $\omega_{2}$ are adjusted to $1.0$ and $5.0$, respectively. This means that we first use the Inter-level loss to initialize the hierarchical coding, followed by incorporating the Cross-level loss to refine the embedding. For tracking loss, we set $\delta=0.99$, $w_{1}=1.0$, $w_{2}=0.5$. For mapping, we set $w_{3}=1.0$, $w_{4}=0.5$, $w_{5}=0.2$, respectively.